Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
berni - 14:58 Tuesday 11 November 2025 (68134) Print this report
Operator Report - Morning shift

ITF found in Science mode.

At 7:00 UTC ITF in Maintenance mode, below the list of the activities communicated in control room:

  • standard tasks performed by the operator:
    • scan and lock of the OMC in single bounce;
    • TCS thermo camera images reference;
    • check of the TCS powers:
  CH [W] OUTER RING [W] INNER RING [W]
WI 0.264 0.295 0.032
NI 0.682 0.57 0.569
  • CRAL building digging work;
  • from 9:30 UTC to 10:30 UTC PSTAb realignment in the AOM (INJ team);
  • SNEB microsprings and DMS warning;
  • Tests of ACS INJ; for this test I set the UTA INJ in PORTATA NOMINALE from 10:13 UTC to 10:49 UTC;

All the activities completed at 11:18 UTC; I manually realigned the cavities, then the CITF and then I adjusted MICH offset. Finally the ITF was in Science mode at 12:40 UTC.

 

 

SUSP

This morning the susp on-call tried to fix the high voltage for F0 OB COIL H2; unfurtunately he discovered a bug in the software that is preventing to restore the situation. At the moment the SUSP team is trying to fix the bug and then they will be able to restore the corrections.

 

 

 

Images attached to this report
Comments to this report:
ballardin - 21:49 Tuesday 11 November 2025 (68145) Print this report

After lunch it was necessary to reboot the olserverwin.virgo.infn.it Windows machine because it was not responding to ping.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×